A company has a manufacturing plant that is producing quality lamps. They find that in order to produce 250 lamps in a month, it will cost $7150. Also, to produce 310 lamps in a month, it will cost $8050. Find an equation in the form y = m x + b, where x is the number of lamps produced in a month and y is the monthly cost to do so.
I figured out that the slope is 15 but I'm not sure how to calculate "b". Can someone please help me?

Answer by macston(5194)

y=mx+b You have x and y given, plug them in with your calculated m and solve.
If m=15 is correct:
y=15x+b Using 250 and $7150:
$7150=15(250)+b Subtract 15(250) from each side.
$7150-15(250)=b
3400=b
CHECK with the other:
y=15x+b Using 310 and $8050
$8050=15(310)+b
$8050-15(310)=b
3400=b
ANSWER: The equation is y=15x+3400.